OpenStudy (anonymous):

How many distant natural numbers are in the solution set of {x:|2x+6|≥4} and (intersection) {x: |3x+20|≤29}

OpenStudy (anonymous):

*distinct

OpenStudy (anonymous):

The first absolute equation breaks down to \[x\ge -1\;or\; x\le -5\] The second ab.eq. becomes \[x\le 3\; and\; x\ge \frac{-49}{3}\] Draw these both on a number line and count.

OpenStudy (anonymous):

wow thanks!